# Simulating and Sampling from Quantum Circuits with 2D Tensor Networks

### v1

@@ Abstract
Classical simulations of quantum circuits play a vital role in the development of quantum computers and for taking the temperature of the field. Here, we classically simulate various physically-motivated circuits using 2D tensor network ansätze for the many-body wavefunction which match the geometry of the underlying quantum processor. We then employ a generalized version of the boundary Matrix Product State contraction algorithm to controllably generate samples from the resultant tensor network states. Our approach allows us to systematically converge both the quality of the final state and the samples drawn from it to the true distribution defined by the circuit. With these methods, we simulate the largest local unitary Jastrow ansatz circuit taken from recent IBM experiments to numerical precision. We also study a domain-wall quench in a two-dimensional discrete-time Heisenberg model on large heavy-hex and rotated square lattices, which reflect IBM’s and Google’s latest quantum processors respectively. We observe a rapid buildup of complex loop correlations on the Google Willow geometry which significantly impact the local properties of the system. Meanwhile, we find loop correlations build up extremely slowly on heavy-hex processors and have almost negligible impact on the local properties of the system, even at large circuit depths. Our results underscore the role the geometry of the quantum processor plays in classical simulability.
@@ I Introduction
Figure 1: Our tensor network topologies and their simulation errors. a) The tensor network states (TNS) for simulating the N2 and 4FE-4S LUCJ circuits according to Ref. [12]. Vertex colors indicate quantum sub-registers which encode the different spin states for the considered electronic system. b) Approximate infidelity (c.f. $1-f$ following Eq. (2)) of the final TNS as a function of the maximum bond dimension $\chi$ in our simulations. We also sketch the circuit structures, which consist of single-site gates and XX+YY rotations within the quantum registers, and CP gates between them. c) The TNS used for simulating discrete-time dynamics under the Heisenberg Hamiltonian on the heavy-hex or Willow processor topologies. Vertex colors indicate the regions of the model initialized in the 0- or the 1-state. d) Approximate infidelity of the final TNS after simulating $20$ layers on the heavy-hex topology, and $7$ or $15$ on the Willow topology. Further details of the simulations can be found in Sec. III.
@@
Here we demonstrate the simulation and sampling of quantum circuits with 2D tensor networks in a controllable and verifiable manner. Our corresponding open-source software [16] can be used to perform robust tensor network simulations of circuits realized over any planar quantum processor. We consider two different circuits: i) the local unitary cluster Jastrow ansatz (LUCJ) circuits employed in Ref. [12] and simulated on cutouts of IBM’s latest processors with 52 and 72 qubits, and ii) the discrete-time dynamics of a domain wall quench of a 2D Heisenberg model on both IBM’s heavy-hex architecture with 164 qubits and the latest Willow processor by Google with 105 qubits [17].
@@ II Methods
with the corresponding infidelity $1-f$. Whilst this quantity is an approximation of the overall infidelity, in practical experience it is often a reliable error metric for the accuracy of the simulation in terms of overall state fidelity [2, 14].
@@
As a consequence, we have the automated ability to systematically contract any planar tensor network in a highly efficient, controllable manner. This then allows us to controllably extract expectation values (including non-local correlators [30]) and, crucial to this work, sample bitstrings $x$ from the tensor networks $|\psi\rangle$ illustrated in Fig. 1 via an implementation of the TNS sampling procedure introduced in Ref. [31] but generalised to arbitrary planar topologies.
@@
[…]
@@
which is the (inverse) log-ratio of the probabilities averaged over the samples drawn from $q(x)$. A KLD of 00 guarantees that the distributions are identical, but small values significantly below $1$ typically indicate high-quality samples.
@@ III Results
Figure 2: Simulation and sampling of local unitary Jastrow circuits. At the top we depict the KLD (see Eq. (II)) of $1000$ samples generated with boundary MPS ov varying different bond dimensions $R$ from a TN state with bond dimension $\chi=600$ after application of the corresponding LUCJ circuit. Annotated times indicate the average time $\bar{t}_{\text{sample}}$ to generate a single bitstring on a Macbook M1, following a pre-computed contraction of the norm network $\langle\psi|\psi\rangle$ for a given $R$. The 72-qubit simulation of the 4Fe-4S molecule exhibits effectively no loop correlations and can be sampled exactly with $R=1$, while the 52-qubit N2 simulation requires at least $R=5$ for near-exact samples. At the bottom we show the distribution of $p(x)/q(x)$ values at sampling dimensions $R=1$ and $R=20$ for the N2 molecule, confirming that at $R\leq 20$ we generate practically exact samples. The inset showcases that samples generated at any $R$ lie within the correct magnetization subspace.
@@
We study the two largest circuits executed in Ref. [12] for the N2 molecule with 52 qubits and the 4Fe-4S molecule with 72 qubits. The executed circuits can be found as part of that work, and consist of particle number preserving rotations - most prominently controlled-phase gates and ${\rm XX}+{\rm YY}$ rotations. In Ref. [12], the ${\rm XX}+{\rm YY}$ gates are transpiled into device-native CNOT gates, with the largest circuit (for the $4$F-$4$s molecule) involving $\sim 3500$ CNOTs. Here, we do not need such transpilation and thus run the same circuit in a format involving $\sim 1800$ ${\rm XX}+{\rm YY}$ rotations. The circuit topologies are sublattices of IBM’s heavy-hex processors with 6 and 4 primitive loops, respectively (see Fig. 1). Here, we show that, with a tensor network graph adapted to the circuit topology, we can simulate the 52-qubit case near-exactly and the 72-qubit case to numerical precision with resources available on a Macbook M1. Note that Ref. [12] reports 58 and 77-qubit simulations, where ancilla qubits had to be used to fit the problem into the heavy hex topology of the quantum device. Here such ancilla are not necessary and can be viewed as just the presence of redundant identity matrices on the bonds of the tensor network (see white nodes in Fig. 1).
@@
Figure 3: Simulation and sampling of the quench dynamics of a domain wall for the Heisenberg model on heavy-hex and Willow topologies. At the top we depict the KLD (see Eq. (II)) of $1000$ samples generated with boundary MPS with indicated bond dimensions $R$. Annotated times indicate the average time $\bar{t}_{\text{sample}}$ to generate a single bitstring on a Macbook M1, following a pre-computed contraction of the norm network $\langle\psi|\psi\rangle$ for a given $R$. At the bottom we show the histogram of the $p(x)/q(x)$ probability ratios at $R=1$ and $R=20$ for the more challenging $L=15$ Willow case. We find that at $R=1$ the probabilities of the samples are often off by dozens of orders of magnitude, which is why the bars within the shown limits are barely visible. In contrast, at $R=20$ we generate higher-quality samples with more more concentrated probability ratios. The inset shows the percentage of samples that lie within the correct magnetization subspace, which is over $96\%$ in all cases at $R=3$.
@@
Figure 4: Expectation values via direct contraction and sampling for the Heisenberg model. We show Pauli $Z$ expectation values for the quench dynamics of a domain wall of the Heisenberg model on sites 51 to 57, which are located the centers of the 164-qubit heavy-hex and 105-qubit Willow topologies, as a function of the boundary MPS dimension $R$. The expectation values are estimated both directly through boundary MPS contraction of $\langle\psi|\psi\rangle$ and via $1000$ samples generated at the designated boundary MPS dimension (1 std error of the mean is shaded). It becomes clear that the Willow topology generates significantly stronger loop correlations than the heavy-hex topology, even at a third of the circuit depth. Interestingly, at sample KLD values of approximately 2 (c.f., Fig. 3), the samples can accurately recover local expectation values while the full distribution still differs from the true one.
@@
Whilst the sample KLD provides an accurate indicator of the quality of our samples, due to its global nature it can be an overly conservative estimate of accuracy when the desired measurement outcomes of the state are low-weight observables such as one or two-site expectation values. In Fig. 4 we showcase the local $Z_{i}$ expectation values for sites close to the center of both topologies — computed both from the sample distributions illustrated in Fig. 7 and from direct computation using MPS messages to contract the $\langle\psi|Z_{i}|\psi\rangle$ and $\langle\psi|\psi\rangle$. The heavy-hex topology shows an almost complete insensitivity to the single-site expectation with the boundary MPS dimension both when sampling and directly computing an observable. The Willow topology is notably different, even at shallower circuit depths. At $7$ layers, expectation values can be converged with our MPS algorithms but require dimensions $R$ on the order of the bond dimension $\chi$ of the state. Meanwhile, at $15$ layers it becomes very difficult to converge the single-site expectation value with either direct computation or sample-based computation. Expectation values derived alternate boundary MPS contraction schemes, as well as BP loop corrections [36, 30] — another recent method for measuring expectation values beyond the BP approximation — are shown in the Appendix in Fig. 6.
@@
To study the loop correlations present in these TNs more directly, we compute the first-order approximation to the BP error in both setups (see Eq. (3)) as a function of the number of Trotter layers. This error can be seen as quantifying the strength of loop correlations in the TNS. The results are shown in Fig. 5. We observe a drastically larger BP error (many orders of magnitude) for the Willow square-lattice versus the heavy-hexagonal lattice. The loops in the heavy-hex lattice are three times as large and thus generically one could expect $\varepsilon({\rm Heavy\ Hex})\sim\varepsilon^{3}({\rm Willow})$ ($0\leq\varepsilon\leq 1$) based on the exponential scaling of the eigenvalue gap of increasingly long sequences of matrices. The difference we see in Fig. 5, however, appears to go beyond even this, and is reinforced by the remarkably accurate BP results numerically observed in Fig. 4 and in Ref. [13] for Ising model dynamics on large heavy-hex geometries. These results point to some level of “loop interference” in large lattice systems which compounds with the increased loop sizes relative to the Willow topology. A theoretical explanation appears urgently needed to this phenomenon.
@@ IV Conclusion
By applying these techniques to the local unitary Jastrow ansatz (LUCJ) circuits introduced in Ref. [12], we generated samples which are drawn, to numerical precision, from the exact underlying distribution. We also showed the generality of our methods, simulating, on moderate timescales, the highly-entangling quench dynamics of a domain wall in the two-dimensional Heisenberg model on both IBM’s heavy-hex processor architecture and Google’s latest Willow processor.
@@
Our results here highlight how geometry plays a crucial role in the complexity of quantum circuits and provide a state-of-the-art framework — with corresponding open-source code [16, 27]— for simulating quantum circuits with planar tensor networks. We hope that these tensor networks and the underlying classical simulation methods in general become more widely used by those at the forefront of developing quantum devices and their applications.
@@ Acknowledgements
Effect of contraction order in the Willow geometry
Figure 6: Boundary MPS expectation values from diagonal contraction and comparison to loop corrections. We compare to the data presented in Fig. 4, where the boundary MPS contractions are not performed column-wise (c.f. Fig. 1) but diagonally. For the same MPS dimension $R$, the results appear more accurate, but each $R$ is also more costly due to a higher typical coordination number. The loop correction expectation values are performed by contracting loops with 0 (BP), 4, or 7 edges exactly. Especially for the 15-layer case, it is not clear whether they are more accurate than the boundary MPS results.
The boundary MPS contractions for the expectation values and sampling in Fig. 4 were performed with a column-wise partitioning of the TN states. In Fig. 6, we instead show results attained by a diagonal contraction of the Willow layout as it is displayed in Fig. 1. The diagonal contract scheme more closely resembles the conventional approach to contracting square lattices, i.e., from one of the sides instead of the corner. One effect is that the partition boundary MPS then have approximately equal lengths and there is less of them, which may improve the MPS-MPO fitting procedure. At the same time, the coordination number of the contractions is now bigger, which makes every contraction more expensive at the same bond dimensions.
In Fig. 6 we confirm this understanding, showing that at the same boundary MPS dimension $R$ the expectation values are more stable (particularly at low $R$ and in the $7$ layer case), but the runtimes of these computations are higher by several factors. As an external source of validation, we also report the BP loop correction results, including exact contraction of loops with 0 (BP), 4, and 7 edges. While these can also not be considered converged, they hint at a region within which the true expectation values are likely to lie.
With our flexible open-source code, this case study opens up the question of which MPS contraction order is best. For the $L=7$ case, we were able to push the contraction of the $\chi=10$ TN state to $R=100$, which gives exact results with expectation values that have long converged. While for the same $R$, the diagonal contractions may be of higher quality, the runtimes and memory requirements are substantially elevated and do currently not allow us to reach $R=100$. It stands to reason that custom partitionings that depend on the local bond dimensions or coordination numbers can be discovered, enabled by our flexible boundary MPS code.
@@
Figure 7: Sampling from planar Tensor Network States. a) Planar tensor network representation of a wavefunction $|\psi\rangle$ with bond dimension $\chi$. A heavy-hex topology is chosen for illustrative purposes. The tensors in the network are grouped into partitions $b=1,2\ldots N_{b}$ with a column-based partition illustrated here. b) Norm network $\langle\psi|\psi\rangle$ with the individual nodes formed from uncontracted pairs of tensors in $|\psi\rangle$. c) MPS-MPO fitting procedures that form the workhorse of the sampling procedure. The contraction of incident MPS with either the MPO defined by the two-layer partition $\psi^{\dagger}_{b}\psi_{b}$ or the MPO defined by the single-layer partition $x_{b}\cdot\psi_{b}$ are approximated with an outgoing MPS via a variational one-site fitting procedure that works on MPOs of arbitrary structure. Other procedures can be used, but the one-site fitting procedure scales most favorably in bond dimension $\chi$. d) Procedure for generating a single bitstring $x$ assuming the pre-computation of the set of MPS $\{M_{N_{b}\rightarrow N_{b}-1},\ldots M_{3\rightarrow 2},M_{2\rightarrow 1}\}$ via the MPS-MPO fitting procedure on the two-layer tensor network state. The first partition $\psi^{\dagger}_{1}\psi_{1}$ is sampled with the MPS $M_{2\rightarrow 1}$ incident. This is done by contracting the structure from top to bottom and sequentially splitting open the bonds connecting the bra and ket tensors to form the one-site reduced density matrix and sample from it, conditioned on the already sampled sites above it. The structure $\langle x_{1}|\psi_{1}\rangle$ is then fit to an MPS $m_{1\rightarrow 2}$ and the second partition is sampled. The structure $m_{1\rightarrow 2}\cdot x_{2}\cdot\psi_{2}$ is then fit to an MPS $m_{2\rightarrow 3}$ and the next partition sampled. This is repeated until all partitions are sampled, yielding a bitstring $x$ from the distribution $q(x)$ defined by the dimension $R_{n}$ and $R_{x}$ chosen for the MPS $M_{i+1\rightarrow i}$ and $m_{i\rightarrow i+1}$ respectively. Computation of $p(x)=|\langle x|\psi\rangle|^{2}$ can be done either by selecting a sufficiently large $R_{x}$ or a separate contraction of the network $\langle x|\psi\rangle$. A legend is included to show the different tensors which appear and the dimension of their respective indices.
@@
Consider a planar tensor network $|\psi\rangle$ which we wish to draw samples $x$ from. In Fig. 7 we illustrate this with the example of a network with heavy-hex topology. The tensors of the network are first grouped into partitions $b=1,2,\dots N_{b}$ such that the topology of the network following this partitioning is a single line with edges between sequential partitions. This is most naturally achieved by partitioning the network by its rows or columns (we show the choice of column partitions in Fig. 7), although other choices are possible (see, e.g., Fig. 6). We define the subset of tensors in a given partition with $\psi_{b}$. We also define the norm network $\langle\psi|\psi\rangle$ and an identical partitioning such that the partitions $T_{b}=\psi^{\dagger}_{b}\psi_{b}$ are formed from the tensors in $\psi_{b}$ and their conjugates. Those $T_{b}$ are generally MPOs, apart from the first and last partition at the boundaries, where they are MPS.

### v2

@@ Abstract
Classical simulations of quantum circuits play a vital role in the development of quantum computers and for taking the temperature of the field. Here, we classically simulate various physically-motivated circuits using 2D tensor network ansätze for the many-body wavefunction which match the geometry of the underlying quantum processor. We then employ a generalized version of the boundary Matrix Product State contraction algorithm to controllably generate samples from the resultant tensor network states. Our approach allows us to systematically converge both the quality of the final state and the samples drawn from it to the true distribution defined by the circuit, with GPU hardware providing us with significant speedups over CPU hardware. With these methods, we simulate the largest local unitary Jastrow ansatz circuit taken from recent IBM experiments to numerical precision. We also study a domain-wall quench in a two-dimensional discrete-time Heisenberg model on large heavy-hex and rotated square lattices, which reflect IBM’s and Google’s latest quantum processors respectively. We observe a rapid buildup of complex loop correlations on the Google Willow geometry which significantly impact the local properties of the system. Meanwhile, we find loop correlations build up extremely slowly on heavy-hex processors and have almost negligible impact on the local properties of the system, even at large circuit depths. Our results underscore the role the geometry of the quantum processor plays in classical simulability.
@@ I Introduction
Figure 1: Our tensor network topologies and their simulation errors. a) The tensor network states (TNS) for simulating the N2 and 4FE-4S LUCJ circuits according to Ref. [12]. Vertex colors indicate quantum sub-registers which encode the different spin states for the considered electronic system. b) Approximate infidelity (c.f. 1f1-f following Eq. (2)) of the final TNS as a function of the maximum bond dimension $\chi$ in our simulations. We also sketch the circuit structures, which consist of single-site gates and XX+$Y$ rotations within the quantum registers, and CP gates between them. c) The TNS used for simulating discrete-time dynamics under the Heisenberg Hamiltonian on the heavy-hex or Willow processor topologies. Vertex colors indicate the regions of the model initialized in the 0- or the 1-state. d) Approximate infidelity of the final TNS after simulating 2020 layers on the heavy-hex topology, and 77 or 1515 on the Willow topology. Further details of the simulations can be found in Sec. III.
@@
Here we demonstrate the simulation and sampling of quantum circuits with 2D tensor networks in a controllable and verifiable manner. Our corresponding open-source software [16] can be used to perform robust tensor network simulations of circuits realized over any planar quantum processor, with GPU support enabling these samples and expectation values to be obtained with unprecedented speed. We consider two different circuits: i) the local unitary cluster Jastrow ansatz (LUCJ) circuits employed in Ref. [12] and simulated on cutouts of IBM’s latest processors with 52 and 72 qubits, and ii) the discrete-time dynamics of a domain wall quench of a 2D Heisenberg model on both IBM’s heavy-hex architecture with 164 qubits and the latest Willow processor by Google with 105 qubits [17].
@@ II Methods
with the corresponding infidelity 1f1-f. Whilst this quantity is an approximation of the overall infidelity, in practical experience it is often a reliable error metric for the accuracy of the simulation in terms of overall state fidelity [2, 14].
@@
As a consequence, we have the automated ability to systematically contract any planar tensor network in a highly efficient, controllable manner. This then allows us to controllably extract expectation values (including non-local correlators [30]) and, crucial to this work, sample bitstrings $x$ from the tensor networks $|\psi\rangle$ illustrated in Fig. 1 via an implementation of the TNS sampling procedure introduced in Ref. [31] but generalised to arbitrary planar topologies. Moreover, as our boundary MPS approach is dominated by tensor contractions and QR decompositions, significant speedups are observed when using GPU hardware — which we will show and exploit explicitly in this work to rapidly and accurately contract 2D tensor networks of very high bond dimension.
@@
$q(x)$ : $:\text{ the sampled distribution.}$
$p(x)$ : $:\text{ the actual distribution }|\langle x|\psi\rangle|^{2}$
 $\text{\hskip 8.53581pt defined by the TNS}.$
@@
which is the (inverse) log-ratio of the probabilities averaged over the samples drawn from $q(x)$. A KLD of 0 guarantees that the distributions are identical, but small values significantly below $1$ typically indicate high-quality samples.
@@ III Results
Figure 2: Simulation and sampling of local unitary Jastrow circuits. At the top we depict the KLD (see Eq. (II)) of $1000$ samples generated with boundary MPS ov varying different bond dimensions $R$ from a TN state with bond dimension $\chi=600$ after application of the corresponding LUCJ circuit. Annotated times indicate the average time $t_{\text{sample}}$ to generate a single bitstring using either an Intel Xeon Gold (CPU) or an Nvidia RTX A6000 (GPU), following a pre-computed contraction of the norm network $\langle\psi|\psi\rangle$ for a given $R$. The 72-qubit simulation of the 4Fe-4S molecule exhibits effectively no loop correlations and can be sampled exactly with $R=1$, while the 52-qubit N2 simulation requires at least $R=5$ for near-exact samples. At the bottom we show the distribution of $p(x)/q(x)$ values at sampling dimensions $R=1$ and $R=20$ for the N2 molecule, confirming that at $R\leq 20$ we generate practically exact samples. The inset showcases that samples generated at any $R$ lie within the correct magnetization subspace.
@@
We study the two largest circuits executed in Ref. [12] for the N2 molecule with 52 qubits and the 4Fe-4S molecule with 72 qubits. The executed circuits can be found as part of that work, and consist of particle number preserving rotations - most prominently controlled-phase gates and ${\rm XX}+{\rm YY}$ rotations. In Ref. [12], the ${\rm XX}+{\rm YY}$ gates are transpiled into device-native CNOT gates, with the largest circuit (for the 44F-44s molecule) involving $\sim 3500$ CNOTs. Here, we do not need such transpilation and thus run the same circuit in a format involving $\sim 1800$ ${\rm XX}+{\rm YY}$ rotations. The circuit topologies are sublattices of IBM’s heavy-hex processors with 6 and 4 primitive loops, respectively (see Fig. 1). Here, we show that, with a tensor network graph adapted to the circuit topology, we can simulate the 52-qubit case near-exactly and the 72-qubit case to numerical precision, with a time-to-sample far below one second on a single GPU. Note that Ref. [12] reports 58 and 77-qubit simulations, where ancilla qubits had to be used to fit the problem into the heavy hex topology of the quantum device. Here such ancilla are not necessary and can be viewed as just the presence of redundant identity matrices on the bonds of the tensor network (see white nodes in Fig. 1).
@@
Figure 3: Simulation and sampling of the quench dynamics of a domain wall for the Heisenberg model on heavy-hex and Willow topologies. At the top we depict the KLD (see Eq. (II)) of $1000$ samples generated with boundary MPS with indicated bond dimensions $R$. Annotated times indicate the average time $t_{\text{sample}}$ to generate a single bitstring using either an Intel Xeon Gold (CPU) or an Nvidia RTX A6000 (GPU). At the bottom we show the histogram of the $p(x)/q(x)$ probability ratios at $R=1$ and $R=20$ for the more challenging $L=15$ Willow case. We find that at $R=1$ the probabilities of the samples are often off by dozens of orders of magnitude, which is why the bars within the shown limits are barely visible. In contrast, at $R=20$ we generate higher-quality samples with more more concentrated probability ratios. The inset shows the percentage of samples that lie within the correct magnetization subspace, which is over $96\%$ in all cases at $R=3$.
@@
Figure 4: Expectation values via direct contraction and sampling for the Heisenberg model. We show Pauli $Z$ expectation values for the quench dynamics of a domain wall of the Heisenberg model on sites 51 to 57, which are located the centers of the 164-qubit heavy-hex and 105-qubit Willow topologies, as a function of the boundary MPS dimension $R$. The expectation values are estimated both directly through boundary MPS contraction of $\langle\psi|\psi\rangle$ and via 10001000 samples generated at the designated boundary MPS dimension (1 std error of the mean is shaded). It becomes clear that the Willow topology generates significantly stronger loop correlations than the heavy-hex topology, even at a third of the circuit depth. Interestingly, at sample KLD values of approximately 2 (c.f., Fig. 3), the samples can accurately recover local expectation values while the full distribution still differs from the true one.
@@
Whilst the sample KLD provides an accurate indicator of the quality of our samples, due to its global nature it can be an overly conservative estimate of accuracy when the desired measurement outcomes of the state are low-weight observables such as one or two-site expectation values. In Fig. 4 we showcase the local $Z_{i}$ expectation values for sites close to the center of both topologies — computed both from the sample distributions illustrated in Fig. 7 and from direct computation using MPS messages to contract the $\langle\psi|Z_{i}|\psi\rangle$ and $\langle\psi|\psi\rangle$. The heavy-hex topology shows an almost complete insensitivity to the single-site expectation with the boundary MPS dimension both when sampling and directly computing an observable. The Willow topology is notably different, even at shallower circuit depths. At $7$ layers, expectation values can be converged with our MPS algorithms but require dimensions $R$ on the order of the bond dimension $\chi$ of the state. Meanwhile, at $15$ layers it becomes much more computationally expensive to converge the single-site expectation value with either direct computation or sample-based computation.
At this depth, we find we have to go to MPS dimensions $R\sim 75$ to obtain convergence in local expectation values when contracting the norm of this 2D tensor network $|\psi\rangle$ which has bond dimension $\chi=20$. These results thus necessitated the accurate contraction of a 2D PEPS of very high bond dimension and we achieve this here because, as we systematically increase boundary MPS rank, the workload of our fitting method is increasingly dominated by tensor contractions and allows us to leverage GPU hardware to its fullest extent. We show a comparison of the relevant walltimes in Fig. 5 on both CPU and GPU, realizing a speedup factor of over $35$ for GPU hardware both when directly contracting $\langle\psi|\psi\rangle$ with our boundary MPS approach and when generating individual samples via boundary MPS contraction.
Figure 5: Walltimes for contracting and sampling from a 2D Willow tensor network. The 105-qubit tensor network state $|\psi\rangle$ with $\chi=20$ is the same as in Fig. 4, i.e that obtained after 1515 layers of the propagator in Eq. (8) with $\delta tJ=0.1$. At the top is the walltime to contract the norm network $\langle\psi|\psi\rangle$ with boundary MPS of dimension $R$. At the bottom is the average time to generate a sample $x$ from $|\psi\rangle$ using a boundary MPS of dimension $R$, following a pre-computed contraction of the norm network $\langle\psi|\psi\rangle$. CPU hardware corresponds to a multi-threaded Intel Xeon 6244 Gold CPU whilst GPU hardware corresponds to an Nvidia RTX A6000. All calculations are in 32-bit floating point precision.
@@
We now study the loop correlations present in these TNs. When these are large, they necessitate the aforementioned contraction with a large boundary MPS dimension $R$. Here, we compute the first-order approximation to the BP error in both setups (see Eq. (3)) as a function of the number of Trotter layers. This error can be seen as quantifying the strength of loop correlations in the TNS. The results are shown in Fig. 6. We observe a drastically larger BP error (many orders of magnitude) for the Willow square-lattice versus the heavy-hexagonal lattice. The loops in the heavy-hex lattice are three times as large and thus generically one could expect $\varepsilon({\rm Heavy\ Hex})\sim\varepsilon^{3}({\rm Willow})$ ($0\leq\varepsilon\leq 1$) based on the exponential scaling of the eigenvalue gap of increasingly long sequences of matrices. The difference we see in Fig. 6, however, appears to go beyond even this, and is reinforced by the remarkably accurate BP results numerically observed in Fig. 4 and in Ref. [13] for Ising model dynamics on large heavy-hex geometries. These results point to some level of “loop interference” in large lattice systems which compounds with the increased loop sizes relative to the Willow topology. A theoretical explanation appears urgently needed to this phenomenon.
@@ IV Conclusion
By applying these techniques to the local unitary Jastrow ansatz (LUCJ) circuits introduced in Ref. [12], we generated samples which are drawn, to numerical precision, from the exact underlying distribution. We also showed the generality of our methods, simulating, on moderate timescales, the highly-entangling quench dynamics of a domain wall in the two-dimensional Heisenberg model on both IBM’s heavy-hex processor architecture and Google’s latest Willow processor. We exploited the potential for GPU speedup latent in these contraction methods, to accurately sample and contract the norm of 2D Tensor Networks of very large bond dimension.
@@
Our results here highlight how geometry plays a crucial role in the complexity of quantum circuits and provide a state-of-the-art framework — with corresponding open-source code with both CPU and GPU support [16, 27] – for simulating quantum circuits with planar tensor networks. We hope that these tensor networks and the underlying classical simulation methods in general become more widely used by those at the forefront of developing quantum devices and their applications.
@@ Acknowledgements
Figure 7: Sampling from planar Tensor Network States. a) Planar tensor network representation of a wavefunction $|\psi\rangle$ with bond dimension $\chi$. A heavy-hex topology is chosen for illustrative purposes. The tensors in the network are grouped into partitions $b=1,2\ldots N_{b}$ - with a column-based partition illustrated here. b) Norm network $\langle\psi|\psi\rangle$ with the individual nodes formed from uncontracted pairs of tensors in $|\psi\rangle$. c) MPS-MPO fitting procedures that form the workhorse of the sampling procedure. The contraction of incident MPS with either the MPO defined by the two-layer partition $\psi^{\dagger}_{b}\psi_{b}$ or the MPO defined by the single-layer partition $x_{b}\cdot\psi_{b}$ are approximated with an outgoing MPS via a variational one-site fitting procedure that works on MPOs of arbitrary structure. Other procedures can be used, but the one-site fitting procedure scales most favorably in bond dimension $\chi$. d) Procedure for generating a single bitstring $x$ assuming the pre-computation of the set of MPS $\{M_{N_{b}\rightarrow N_{b}-1},\ldots M_{3\rightarrow 2},M_{2\rightarrow 1}\}$ via the MPS-MPO fitting procedure on the two-layer tensor network state. The first partition $\psi^{\dagger}_{1}\psi_{1}$ is sampled with the MPS $M_{2\rightarrow 1}$ incident. This is done by contracting the structure from top to bottom and sequentially splitting open the bonds connecting the bra and ket tensors to form the one-site reduced density matrix and sample from it, conditioned on the already sampled sites above it. The structure $\langle x_{1}|\psi_{1}\rangle$ is then fit to an MPS $m_{1\rightarrow 2}$ and the second partition is sampled. The structure $m_{1\rightarrow 2}\cdot x_{2}\cdot\psi_{2}$ is then fit to an MPS $m_{2\rightarrow 3}$ and the next partition sampled. This is repeated until all partitions are sampled, yielding a bitstring $x$ from the distribution $q(x)$ defined by the dimension $R_{n}$ and $R_{x}$ chosen for the MPS $M_{i+1\rightarrow i}$ and $m_{i\rightarrow i+1}$ respectively. Computation of $p(x)=|\langle x|\psi\rangle|^{2}$ can be done either by selecting a sufficiently large $R_{x}$ or a separate contraction of the network $\langle x|\psi\rangle$. A legend is included to show the different tensors which appear and the dimension of their respective indices.
@@
Consider a planar tensor network $|\psi\rangle$ which we wish to draw samples $x$ from. In Fig. 7 we illustrate this with the example of a network with heavy-hex topology. The tensors of the network are first grouped into partitions $b=1,2,\dots N_{b}$ such that the topology of the network following this partitioning is a single line with edges between sequential partitions. This is most naturally achieved by partitioning the network by its rows or columns (we show the choice of column partitions in Fig. 7), although other choices are possible. We define the subset of tensors in a given partition with $\psi_{b}$. We also define the norm network $\langle\psi|\psi\rangle$ and an identical partitioning such that the partitions $T_{b}=\psi^{\dagger}_{b}\psi_{b}$ are formed from the tensors in $\psi_{b}$ and their conjugates. Those $T_{b}$ are generally MPOs, apart from the first and last partition at the boundaries, where they are MPS.